An organic compound is composed of Carbon, hydrogen, oxygen, and Chlronie.

  • One mole of the compound reacts completely with one mole of ethanol. If 1.57g of the compound required 0.92g of ethanol. Find the molecular weight and give the molecular formular of the compound
  • The compound hydrolyses readily in water. Give its structural formula, name and write the equation for the reaction.
1
Expert's answer
2021-03-09T04:59:12-0500

n(ethanol) = 0.92 / 46 = 0.02 (mol)

M(substance) = 1.57 / 0.02 = 78.5 (g/mol)

78.5 - 35.5 - 16 = 27 (it is C2H3)

The formula of compound is CH3COCl, it is called acetyl chloride

CH3COCl + H2O = CH3COOH + HCl